Barium ion with a 2+ charge travel at a speed 800 m/s is moving perpendicular to the magnetic field with a magnitude of 2.0 T. Determine the magnitude of magnetic force acting on Ba^(2+).

1
Expert's answer
2022-05-20T13:19:48-0400
"F=qvB\sin\theta\\
=2*1.6*10^{-19}*800*2.0\sin90^\circ\\
=5.12*10^{-16}\:\rm N"